OpenStudy (anonymous):

4=2/3x+9-1/3x pls help me solve this thank you

OpenStudy (anonymous):

4=1/3x+9 -5=1/3x -15x=1 x=-1/15

OpenStudy (anonymous):

the first step is because 2/3x-1/3x=1/3x (like denominators)

OpenStudy (anonymous):

4=2/3x+9-1/3x ===> (1/3)x =4-9 ===> (1/3)x=-5 ==>x=-15